# GlitchMiner: Mining Glitch Tokens in Large Language Models via Gradient-based Discrete Optimization

### v4

@@ Abstract
Glitch tokens in Large Language Models (LLMs) can trigger unpredictable behaviors, threatening model reliability and safety. Existing detection methods rely on predefined patterns, limiting their adaptability across diverse LLM architectures. We propose GlitchMiner, a gradient-based discrete optimization framework that efficiently identifies glitch tokens by introducing entropy as a measure of prediction uncertainty and employing a local search strategy to explore the token space. Experiments across multiple LLM architectures demonstrate that GlitchMiner outperforms existing methods in detection accuracy and adaptability, achieving over 10% average efficiency improvement. This method enhances vulnerability assessment in LLMs, contributing to the development of more robust and reliable applications. Code is available at https://github.com/wooozihui/GlitchMiner.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized natural language processing, enabling breakthrough applications in fields ranging from code generation LLM_code1 ; LLM_code2 ; LLM_code3 to medical research medical1 ; medical2 and education education1 ; education2 . However, with the growing reliance on these models in critical domains, a subtle yet significant vulnerability has emerged: glitch tokens geiping2024coercing . These anomalous tokens can severely disrupt an LLM’s output, leading to repetitive errors, nonsensical responses, or even harmful content. Such disruptions are particularly concerning in high-stakes applications where model reliability is crucial.
Recent research has made progress in detecting glitch tokens. Methods such as Magikarp Magikarp and GlitchHunter GlitchHunter focus on abnormal embedding patterns and clustering properties. However, these approaches are limited by their dependence on predefined patterns, restricting their adaptability across different LLM architectures and potentially missing novel forms of glitch tokens that deviate from known patterns.
In this paper, we introduce GlitchMiner, a novel framework for glitch token detection that overcomes the limitations of pattern-based heuristics by employing gradient-based discrete optimization to directly analyze a model’s prediction behavior. Central to our method is the use of entropy as an exploratory metric, guiding the search through the token space to identify regions indicative of unexpected model behavior.
The key innovations of GlitchMiner include:
  • Entropy-Based Optimization: By introducing entropy as a novel objective for glitch token detection, we systematically identify tokens that induce high uncertainty in model predictions. This approach enables detection of a broader range of glitch tokens, extending beyond traditional pattern- or embedding-based methods.
  • Gradient-Guided Local Search: Combining first-order Taylor approximation with a local search strategy, GlitchMiner efficiently explores the vast token space. This significantly enhances both the efficiency and accuracy of glitch token detection, allowing thorough exploration while maintaining computational feasibility.
  • Adaptive Cross-Architecture Detection: GlitchMiner is designed to adapt to the unique characteristics of various LLM architectures without relying on model-specific assumptions, enabling effective glitch token detection across a wide range of models.
We conducted extensive experiments across 10 popular LLM architectures, covering models from the Llama, Qwen, Gemma, Phi, and Mistral families. The results show that GlitchMiner surpasses existing methods in detection efficiency in 90% of cases. Specifically, GlitchMiner achieves an average Detected@1000 score of 612.0 and a Detected@2000 score of 980.1, with respective improvements of 10.5% and 10.7% over state-of-the-art methods. These gains enhance the detection of potentially harmful tokens, improving both accuracy and query efficiency. Such advancements are essential for LLM reliability, especially in applications where undetected glitch tokens could have serious consequences.
The remainder of this paper is structured as follows: Section 2 provides a review of related work in glitch token detection and gradient-based optimization techniques. Section 3 details the design principles and implementation of GlitchMiner. Section 4 presents our experimental results and analysis, demonstrating the effectiveness of our approach across various LLM architectures. Finally, Section 5 concludes with a discussion of the implications of our work and outlines promising directions for future research.
@@ 2 Background and Related Work
2.1 What is a Glitch Token?
Glitch tokens, a concept that gained attention following the discovery of the "SolidGoldMagikarp" phenomenon by Rumbelow et al. solidgoldmagikarp , are anomalous tokens in LLMs that can trigger unexpected and often erroneous behaviors when processed by the model. These tokens typically result from irregularities in the training process, such as underrepresentation in the training data or inconsistencies in tokenization Magikarp ; geiping2024coercing . Glitch tokens have been found to exhibit several key characteristics:
  • They often have abnormal embedding patterns, such as unusually small $\ell_{2}$ norms or atypical positions in the embedding space Magikarp ; solidgoldmagikarp .
  • When input to an LLM, they can cause the model to produce repetitive, nonsensical, or completely unrelated outputs solidgoldmagikarp ; GlitchHunter .
  • They may lead to significant deviations in the model’s internal activations compared to normal tokens GlitchProber .
  • Glitch tokens can sometimes bypass content filters or trigger unexpected model behaviors, potentially compromising the safety and reliability of LLMs geiping2024coercing .
The study of glitch tokens is crucial for understanding and mitigating potential vulnerabilities in LLMs, as these tokens can potentially be exploited for adversarial attacks or lead to unintended model behaviors in critical applications geiping2024coercing .
@@ 2.2 Glitch Token Detection
The detection of glitch tokens in LLMs has become an increasingly important area of research, given their potential to disrupt model performance and reliability. The fundamental approach to identifying these anomalous tokens often relies on repetition tasks, a method popularized by the SolidGoldMagikarp study solidgoldmagikarp . In these tasks, the model is prompted to repeat a given token, with the assumption that normal tokens should be easily reproducible, while glitch tokens often lead to failures in this simple task.
Building upon this basic principle, researchers have developed several sophisticated methods to detect glitch tokens more efficiently and accurately:
Magikarp solidgoldmagikarp pioneered a fast and lightweight approach by analyzing token embeddings. It identifies potentially problematic tokens by examining characteristic patterns in the embedding space, such as unusual $\ell_{2}$ norms. These candidate tokens are then verified using repetition tasks, allowing for quick detection through direct queries to the model.
GlitchHunter GlitchHunter introduced a clustering-based method. This approach is founded on the observation that glitch tokens often cluster together in the embedding space. GlitchHunter constructs a Token Embedding Graph (TEG) to represent token relationships and applies the Leiden algorithm traag2019louvain to identify potential glitch clusters. These clusters are then refined through iterative hypothesis testing to improve detection accuracy.
GlitchProber GlitchProber takes a different approach by analyzing the internal activations within transformer layers, such as attention heads and hidden states. It reduces the dimensionality of these activations using PCA PCA and applies SVM classifiers SVM to identify glitch tokens. Additionally, it integrates mitigation mechanisms by modifying neuron activations to minimize the impact of glitch tokens during inference.
While these methods have significantly advanced our ability to detect glitch tokens, they are inherently limited by their reliance on pattern-based heuristics. Magikarp and GlitchHunter focus on specific embedding patterns or clustering behaviors, while GlitchProber examines predefined patterns in activation spaces. This dependence on pattern-based approaches can constrain their adaptability across diverse LLM architectures and potentially overlook novel forms of glitch tokens.
Our proposed method, GlitchMiner, aims to address these limitations by employing gradient-based discrete optimization to directly analyze an LLM’s prediction behavior, potentially offering a more flexible and comprehensive approach to glitch token detection.
@@ 2.3 Gradient-based Discrete Optimization
Gradient-based discrete optimization methods hotflip ; autoprompt ; GCG ; PEZ leverage gradient information to predict how individual tokens impact the loss function. These approaches typically treat the one-hot encoding of tokens or token embeddings as continuous vectors to compute gradients, guiding token replacements for optimization.
HotFlip hotflip uses the one-hot encoding of tokens to compute gradients and selects the token with the largest negative gradient to replace the current token, aiming to minimize the loss. However, it only evaluates one candidate token per iteration, which can lead to suboptimal predictions and reduced accuracy.
AutoPrompt autoprompt improves upon HotFlip by evaluating multiple candidate tokens in each iteration. Instead of relying on gradients from one-hot encodings, it utilizes token embedding gradients for loss estimation, enhancing prediction accuracy by considering a broader range of potential token replacements.
GCG GCG extends HotFlip by incorporating multi-candidate token selection, similar to AutoPrompt, but it still uses the one-hot encoding of tokens to compute gradients for loss estimation. Notably, GCG has been applied to automated jailbreaks shen2023anything in LLMs, efficiently searching for adversarial suffixes.
AutoPrompt and GCG both rely on large batch sampling to mitigate inaccuracies in gradient prediction. We identified that these inaccuracies arise from the inaccuracy of Taylor expansions when input tokens are distant from the original points. This overlooks a fundamental condition of Taylor approximation: its accuracy is highest for points close to the reference point.
@@
3 Method
Algorithm 1 GlitchMiner Framework
1:Input: Token set $\mathcal{T}$, Iteration number $I$, Batch size $B$
2:Output: Glitch token set $\mathcal{G}$
3:# Stage 1: Initialization
4:$\mathcal{T}^{*}\leftarrow\text{Filter}(\mathcal{T})$ Filter out the unnecessary token set $\mathcal{T}^{*}$
5:$\mathcal{G}\leftarrow\emptyset$
6:# Stage 2: Mining
7:for $i\in I$ do
8:     Select a batch of tokens $\mathcal{B}$ from $\mathcal{T}\setminus(\mathcal{T}^{*}\cup\mathcal{G})$, where $|\mathcal{B}|=B$
9:     for each token $t\in\mathcal{B}$ do
10:         if Verify($t$then
11:              Add $t$ to $\mathcal{G}$
12:         else
13:              Add $t$ to $\mathcal{T}^{*}$
14:         end if
15:     end for
16:end for
GlitchMiner Framework. As shown in Algorithm 1, GlitchMiner operates in two main stages: initialization and mining.
In the initialization stage, we filter out tokens that don’t require detection to reduce the search space and improve efficiency, after which the mining stage iteratively selects batches of unverified tokens for verification. Each token is checked: if it is identified as a glitch token, it is added to the discovered set; otherwise, it is marked unnecessary for further detection. This process continues until a stopping condition is met, such as reaching a specified number of iterations or detected glitch tokens.
The following sections detail how GlitchMiner implements its key components:
@@ 3 Method
    Token Filter: Identifying tokens that don’t require detection.
@@
    Glitch Token Verification: Methods to verify if a token is a glitch token.
@@
    Token Selection Strategy: Selecting the most promising glitch token candidates in each iteration.
By exploring these components, we demonstrate how GlitchMiner efficiently implements glitch token mining in practice.
3.1 Token Filter
In the initialization stage of GlitchMiner, we first filter all tokens to remove those that do not need to be detected. Similar to Magikarp , we designed a token filter module that classifies tokens based on their characteristics and filters out those that are unnecessary for detection.
The core idea of the token filter module is to classify each token by decoding and then re-encoding it, ensuring it meets specific classification criteria. Specifically, we prepend a special prefix "«" to each token to maintain consistency during the encoding and decoding process. Then, we filter out tokens that fall into the following categories:
  • SPECIAL: Special tokens, such as [BOS], </s>, etc., which are enclosed in brackets or angle brackets and contain alphabetic characters.
  • UNDECODEABLE: Tokens that cannot be decoded, usually containing illegal characters.
  • UNREACHABLE: Tokens that cannot be restored to their original token ID through the decoding and re-encoding process.
During the classification process, we first decode each token ID to obtain its corresponding string representation. If decoding fails, the token is classified as UNDECODEABLE. Next, we encode the decoded string and check if it can be restored to the original token ID. If it cannot, the token is classified as UNREACHABLE. If it meets the characteristics of a special token, it is classified as SPECIAL.
By filtering out these unnecessary tokens, we retain only the essential ones for subsequent mining, which improves the algorithm’s efficiency and ensures the accuracy of the detection results.
3.2 Glitch Token Verification
Our glitch token verification process consists of two stages: optimization-driven detection and cross-verification.
Template for Optimization. For our primary detection and optimization process, we designed the following task template:
User: Please repeat the string: "«{token}»"
Assistant: Sure, the string is: "«{token}
In this template, the blue text represents the prompt, the green text {token} is the input token being tested, and the red text {token} represents the model’s predicted output. If the predicted token does not match the input token, it is initially classified as a potential glitch token.
This template is specifically designed to support our gradient-based optimization process. Its simplicity allows for clear gradient computations, while the prefilled assistant response ensures a uniform starting point for each token evaluation. Wrapping the token in "«" and "»" symbols minimizes interference from surrounding context. These features enable precise entropy calculations and gradient estimations, which are crucial for our entropy-guided search strategy described in the token selection section.
Cross-Verification. While our primary template is optimized for the search process, relying solely on one template may lead to false positives. To enhance the robustness of our detection, we implement a cross-verification step using two additional templates derived from GlitchHunter and Magikarp. Each potential glitch token identified by our primary template undergoes verification with these additional templates. A token is confirmed as a glitch token only if it fails the repetition task across all templates.
This two-stage approach combines the efficiency of our optimization-driven search with the reliability of multi-template verification, significantly reducing false positives while maintaining the effectiveness of our gradient-based detection method.
3.3 Token Selection
Motivation. Previous methods for glitch token detection often rely on manually observed embedding patterns. However, these approaches can be limited in their ability to adapt to diverse LLM architectures and may overlook novel forms of glitch tokens. To address these limitations, we propose an entropy-based approach. Entropy, as a measure of uncertainty in probability distributions, offers a model-agnostic way to identify tokens that cause unexpected behavior in LLMs. By focusing on tokens that maximize entropy, we can detect glitch tokens that deviate significantly from normal token behavior, regardless of their specific characteristics or the underlying model architecture.
Entropy-Guided Exploration. We define the entropy $H(t)$ for a token $t$ as:
where $\mathbf{h}(t)$ is the context embedding for token $t$, and $P(v\mid\mathbf{h}(t))$ is the model’s predicted probability distribution over the vocabulary $\mathcal{V}$. By maximizing entropy, we aim to find tokens that cause high uncertainty in the model’s predictions, potentially indicating glitch behavior.
Optimization Objective. Our goal is to find a batch of tokens $\mathcal{B}$ that maximizes the total entropy:
Here, $\mathcal{T}_{c}$ represents the current candidate set, defined as $\mathcal{T}_{c}=\mathcal{T}\setminus(\mathcal{T}^{*}\cup\mathcal{G})$, where:
  • $\mathcal{T}$ is the initial set of all tokens in the model’s vocabulary.
  • $\mathcal{T}^{*}$ is the set of tokens that have been filtered out or verified as non-glitch tokens. This includes special tokens, undecoded tokens, and tokens that have been checked and confirmed as normal.
  • $\mathcal{G}$ is the set of identified glitch tokens discovered so far in the process.
By excluding $\mathcal{T}^{*}$ and $\mathcal{G}$ from $\mathcal{T}$, we ensure that our search focuses only on the remaining unverified tokens, improving efficiency and preventing redundant checks.
Local Search Strategy. To efficiently solve this optimization problem while addressing the limitations of global Taylor approximation, we introduce a local search strategy. We begin by selecting an initial token $t_{0}$ and computing its entropy $H(t_{0})$ and gradient $\nabla_{e}H(t_{0})$. We then define a local neighborhood $\mathcal{N}_{K}(t_{0})$ consisting of the $K$ nearest neighbors to $t_{0}$ in the embedding space. This localized approach allows us to refine the search within a region close to $t_{0}$, improving approximation accuracy and ensuring efficient exploration of potential glitch tokens.
Within this local neighborhood, we estimate the entropy of each candidate token $t\in\mathcal{N}_{K}(t_{0})$ using first-order Taylor approximation:
$\hat{H}(t)\approx H(t_{0})+\nabla_{e}H(t_{0})^{T}(e_{t}-e_{t_{0}})$
where $e_{t}$ and $e_{t_{0}}$ are the embedding vectors of tokens $t$ and $t_{0}$, respectively. This approximation allows us to efficiently estimate the entropy change for nearby tokens without the computational cost of exact calculations.
Based on these estimates, we select a batch $\mathcal{B}$ of $B$ tokens with the highest approximated entropy. For this selected batch, we then compute the actual entropy values and choose the token with the highest entropy as the starting point for the next iteration. This iterative process ensures that our search is guided by the most promising candidates at each step.
This local search strategy significantly improves the accuracy of entropy estimation by focusing on tokens close to the reference point, addressing a key limitation of previous gradient-based methods. By balancing exploration of the token space with exploitation of local information, our approach enables more efficient and effective glitch token detection.
The combination of entropy-guided exploration and local search allows our method to adapt to diverse LLM architectures without relying on model-specific assumptions. This results in a more robust and versatile approach to glitch token detection, capable of identifying a wide range of anomalous tokens across different model types.
4 Experiments
4.1 Experimental Setup
Test LLMs. We used a diverse set of LLMs from five different model families to evaluate the performance of our glitch token detection approach. The selected models include Meta’s Llama series llama2 ; llama3_1 , Alibaba’s Qwen models qwen2 ; qwen2_5 , Google’s Gemma models gemma2 , Microsoft’s Phi-3 models phi3 , and Mistral models mistral_7B ; mistral2024nemo . The details are presented in Table 1.
@@ 4.1 Experimental Setup
Evaluation Metrics. To assess the performance of our glitch token detection method, we employ the Detected@N metric. This metric measures the absolute number of true glitch tokens identified within the top $N$ predictions made by the algorithm. It is calculated as:
[…]
For example, Detected@1000 measures the number of true glitch tokens discovered within the top 1000 predicted tokens. This metric directly reflects a method’s effectiveness in identifying glitch tokens under practical constraints, as it captures both the detection capability and query efficiency of the approach. By comparing the number of glitch tokens detected under the same query budget ($N$), we can effectively evaluate different methods’ ability to maximize token discovery while minimizing computational resources. This makes it an ideal metric for real-world applications where efficient detection of glitch tokens is crucial.
Baselines. We compare our proposed glitch token detection method with two state-of-the-art approaches: GlitchHunter GlitchHunter and Magikarp Magikarp . These methods serve as the primary benchmarks for evaluating our approach.
Although GlitchProber GlitchProber is another relevant method, it follows a fundamentally different approach by pre-collecting a subset of glitch tokens to train a classifier, introducing a supervised learning component. In contrast, GlitchMiner, along with GlitchHunter and Magikarp, uses heuristic-based methods to detect glitch tokens without relying on labeled data or additional classifier training. This methodological difference makes a direct comparison less meaningful, so we focus our evaluation on methods that align more closely with our unsupervised approach.
Parameter Settings. In our implementation of GlitchMiner, we use $K$=32 and $B$=8 as the default parameters. These values were chosen based on empirical testing to balance computational efficiency and detection effectiveness. Specifically, $K$=32 defines the size of the local neighborhood considered in each iteration, while $B$=8 determines the batch size for entropy computation. These settings have shown to provide a good trade-off between exploration of the token space and exploitation of local information across various model architectures.
4.2 Main Results
@@
Table 2 displays the performance comparison of GlitchMiner with the state-of-the-art methods GlitchHunter and Magikarp across different LLM architectures using the Detected@N metric. The results highlight GlitchMiner’s strong and versatile detection capabilities.
In the majority of cases, GlitchMiner achieved the highest Detected@2000 score, consistently outperforming both baselines in terms of identified glitch tokens within the top 2000 predictions. This is particularly evident in models such as Llama-2-7B-chat-hf, where GlitchMiner achieved a Detected@2000 of 532, surpassing Magikarp by a notable margin. For the Qwen and other models, GlitchMiner maintained robust performance, demonstrating adaptability across different LLM architectures.

### v5

@@ Abstract
Glitch tokens—inputs that trigger unpredictable or anomalous behavior in Large Language Models (LLMs)—pose significant challenges to model reliability and safety. Existing detection methods primarily rely on heuristic embedding patterns or statistical anomalies within internal representations, limiting their generalizability across different model architectures and potentially missing anomalies that deviate from observed patterns. We introduce GlitchMiner, an behavior-driven framework designed to identify glitch tokens by maximizing predictive entropy. Leveraging a gradient-guided local search strategy, GlitchMiner efficiently explores the discrete token space without relying on model-specific heuristics or large-batch sampling. Extensive experiments across ten LLMs from five major model families demonstrate that GlitchMiner consistently outperforms existing approaches in detection accuracy and query efficiency, providing a generalizable and scalable solution for effective glitch token discovery. Code is available at https://github.com/wooozihui/GlitchMiner.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities and are increasingly deployed in high-stakes domains such as code generation (LLM_code1; LLM_code2; LLM_code3), healthcare (medical1; medical2), and education (education1; education2). As reliance on LLMs grows, ensuring their safety and output reliability becomes imperative.
A particularly concerning threat arises from a class of anomalous inputs known as glitch tokens (solidgoldmagikarp). These tokens can induce LLMs to generate unexpected, erratic, or even policy-violating outputs (geiping2024coercing), undermining trust in their behavior. For example, as illustrated in Figure 1, when prompted to repeat the word “Mediabestanden,” Llama2-7b-chat outputs “hello world”—a semantically unrelated response. This phenomenon indicates that certain tokens can trigger unpredictable model behavior, even under simple and deterministic instructions.
In response, several studies have been proposed to detect glitch tokens (Magikarp; GlitchHunter; GlitchProber). These methods typically rely on heuristic observations or statistical patterns in token embeddings or latent representations, such as small embedding norms or localized embedding clusters. Although effective in certain scenarios, such heuristics have two main limitations. First, the reliance on specific observed statistical features may not generalize well across different architectures or training setups. Second, tokens that do not conform to these empirically observed patterns may remain undetected, potentially limiting the overall coverage and robustness of these detection methods.
Human Please repeat the string: Mediabestanden
LLM String: ”hello world”
Figure 1: An illustrative example of how glitch token causing Llama2-7b-chat to fail a simple repetition task. More examples can be found in Appendix B.
To address these limitations, we propose a behavior-driven perspective. Rather than searching for fixed patterns within the model, we hypothesize that glitch tokens are outliers in the model’s learned distribution, and thus elicit uncertain predictions. This uncertainty can be quantified via the entropy of the model’s next-token distribution—high entropy indicates that the model is indecisive, providing a proxy for behavioral instability.
Building on this insight, we present GlitchMiner, a novel framework that identifies glitch tokens by maximizing output entropy. The core of GlitchMiner is a gradient-guided local search, which efficiently navigates the discrete token space to identify high-entropy candidates. Unlike prior gradient-based discrete optimization strategies that suffer from inaccurate approximations or require large batch sampling (autoprompt; GCG), our method restricts updates to neighboring tokens in the embedding space. This ensures accurate Taylor-based entropy estimates and improves search efficiency.
Our Contributions
are summarized as follows:
  • We propose a behavior-driven framework for glitch token detection, leveraging output entropy as an architecture-agnostic measure of model uncertainty to ensure broad applicability across diverse models.
  • We introduce a gradient-guided local search algorithm that accurately estimates entropy gradients within the discrete token space, eliminating the dependence on large-batch sampling and improving search efficiency.
  • Extensive experiments across 10 LLMs from 5 major model families demonstrate that GlitchMiner consistently surpasses state-of-the-art methods in detection accuracy, establishing it as a robust and effective solution for glitch token detection.
The remainder of the paper is organized as follows: Section 2 reviews related work; Section 3 describes the GlitchMiner methodology; Section 4 presents experiments and ablations; and Section 5 concludes with future directions.
@@ 2 Background and Related Work
2.1 Glitch Token Definition
In practice, the identification of glitch tokens relies on specific behavioral tests. The most common approach, used in prior work (Magikarp; GlitchHunter; GlitchProber), is the repetition task. In such a task, a model is prompted to repeat a given token $t$; a failure to accurately reproduce the token is taken as evidence of it being a glitch. This single-test method, however, is fragile. The outcome can be highly sensitive to the specific wording of the prompt, leading to inconsistent results and potential false positives.
To establish a more robust and replicable definition, we move beyond single-prompt evaluations. We classify a token $t$ as a glitch token only if it fails the repetition task consistently across a diverse set of $m$ prompt templates. This stricter criterion is formalized as follows: given a set of $m$ templates, $h_{i}(\cdot)\}_{i=1}^{m}$, the failure rate for a token $t$ is defined as:
$\text{fail}(t):=\frac{1}{m}\sum_{i=1}^{m}\mathds{1}\left[\arg\max_{v\in\mathcal{V}}\,P(v\mid h_{i}(t))\neq t\right]$ (1)
A token is then identified as a glitch if and only if its failure rate is absolute, i.e., $\text{fail}(t)=1$. This rigorous cross-verification process effectively filters out template-specific artifacts, ensuring that our analysis focuses exclusively on consistently unstable tokens.
@@ 2.2 Glitch Token Detection
A series of methods have been proposed to efficiently rank or localise glitch tokens in large vocabularies:
Magikarp (Magikarp) adopts a lightweight heuristic, screening tokens with atypically small $\ell_{2}$ embedding norms before verifying them via the repetition task.
GlitchHunter (GlitchHunter) observes that glitch tokens tend to cluster in embedding space. It builds a token-embedding graph and applies Leiden clustering (traag2019louvain), followed by iterative hypothesis testing to refine each cluster.
GlitchProber (GlitchProber) shifts the focus from embeddings to internal activations. It projects hidden states and attention outputs with PCA, then trains SVM classifiers to flag activation outliers; a mitigation step masks offending neurons at inference time.
While effective, these approaches share a common limitation: they rely heavily on heuristic observations of embedding or activation patterns. Such reliance may reduce their generalizability across different architectures, and tokens not exhibiting these typical patterns may remain undetected.
In contrast, GlitchMiner employs gradient-guided discrete optimization to maximize output entropy, an architecture-agnostic indicator of predictive uncertainty that facilitates broad generalization across diverse LLMs.
@@ 2.3 Gradient-based Discrete Optimization
Gradient-based discrete optimization methods (hotflip; autoprompt; GCG; PEZ) leverage gradient information to predict how individual tokens impact the loss function. These approaches typically treat the one-hot encoding of tokens or token embeddings as continuous vectors to compute gradients, guiding token replacements for optimization.
HotFlip (hotflip) uses the one-hot encoding of the tokens to calculate the gradients and selects the token with the largest negative gradient to replace the current token, with the goal of minimizing the loss. However, it only evaluates one candidate token per iteration, which can lead to suboptimal predictions and reduced accuracy.
AutoPrompt (autoprompt) improves upon HotFlip by evaluating multiple candidate tokens in each iteration. Instead of relying on gradients from one-hot encodings, it utilizes token embedding gradients for loss estimation, enhancing prediction accuracy by considering a broader range of potential token replacements.
GCG (GCG) extends HotFlip by incorporating multi-candidate token selection, similar to AutoPrompt, but it still uses the one-hot encoding of tokens to compute gradients for loss estimation. Notably, GCG has been applied to automated jailbreaks (shen2023anything) in LLMs, efficiently searching for adversarial suffixes.
AutoPrompt and GCG both rely on large batch sampling to mitigate inaccuracies in gradient-based prediction. We identified that these inaccuracies arise from the inaccuracy of Taylor expansions when input tokens are distant from the original points. This overlooks a fundamental condition of Taylor approximation: its accuracy is highest for points close to the reference point.
@@
3 Methodology
Our goal is to develop a method that can automatically and efficiently discover glitch tokens in any given LLM. To this end, we propose GlitchMiner, a novel framework that formulates glitch token discovery as a behavior-driven optimization problem.
GlitchMiner systematically mines for these tokens through a powerful iterative loop, outlined in Algorithm 1. Each iteration first performs Gradient-Guided Selection to pinpoint tokens with the highest predicted output entropy. These candidates then immediately undergo Robust Verification to determine if they are genuine glitch tokens.
Algorithm 1 The GlitchMiner Pipeline
1:Input: LLM model $f(\cdot)$, Full vocabulary $\mathcal{T}$, Max iterations NN
2:Output: Verified glitch token set $\mathcal{G}$
3:# Initialization
[…]
6:# Iterative Mining Loop
7:for $i$ from 11 to NN do
8:  # 1. Gradient-guided selection of promising candidates
9:  $\mathcal{T}_{c}\leftarrow\mathcal{T}\setminus(\mathcal{T}^{*}\cup\mathcal{G})$ $\triangleright$ Define current candidate set
10:  $\mathcal{B}\leftarrow\text{SelectCandidateBatch}(\mathcal{T}_{c})$ $\triangleright$ See Sec. 3.1
11:  # 2. Robust verification of each candidate
12:  for each token $t\in\mathcal{B}$ do
13:   if VerifyIsGlitch($t$) then $\triangleright$ See Sec. 3.2
14:     $\mathcal{G}\leftarrow\mathcal{G}\cup\{t\}$
15:   end if
16:   $\mathcal{T}^{*}\leftarrow\mathcal{T}^{*}\cup\{t\}$ $\triangleright$ Mark token as processed
17:  end for
18:end for
19:return $\mathcal{G}$
3.1 Gradient-Guided Candidate Selection
The core of GlitchMiner’s discovery engine is the candidate selection process. To ensure the overall mining is efficient, this selection is not random; instead, it employs a local search strategy to intelligently find tokens that maximize model confusion, which we measure using output entropy.
Optimization Objective. To formally quantify a token’s impact on model uncertainty, we calculate its output entropy within a repetition task. We use the following template for this optimization process:
User: Please repeat the string: “{token}”
Assistant: Sure, the string is: “{token}
Figure 2: Visualization of GlitchMiner’s local search process. The pivot token (black, $t_{c}$) serves as the reference point. Its neighbor tokens (orange and red) represent the $K=4$ closest tokens in embedding space. Among these, the candidate batch tokens (red) are the top $B=2$ tokens with the highest approximate entropy values, estimated via first-order Taylor approximation. Tokens outside the neighborhood (gray) are excluded to maintain approximation accuracy and computational efficiency.
Here, the first {token} represents the input token $t$ being evaluated, and the second {token} is the next predicted token by LLM. To ensure that the input token appears directly as the model’s next prediction, we prefill the assistant’s response with the phrase Sure, the string is: “. The entropy is computed based on the model’s next-token distribution $p(v\mid\mathbf{h}(t))$, where $\mathbf{h}(t)$ denotes the prompt generated by inserting $t$ into the repetition template $h$. The entropy is thus:
Our goal is to find a batch of tokens $\mathcal{B}$ from the current set of available tokens $\mathcal{T}_{c}$ that maximizes the total entropy:
where $\mathcal{T}_{c}=\mathcal{T}\setminus(\mathcal{T}^{*}\cup\mathcal{G})$, with $\mathcal{T}^{*}$ being the set of filtered (see Section 3.3) or previously verified non-glitch tokens, and $\mathcal{G}$ being the set of glitch tokens found.
Local Search Strategy. To efficiently solve this optimization problem, we introduce a local search strategy that addresses the limitations of global Taylor approximations. This process begins with an initial pivot token $t_{c}$ and iteratively refines the search. In each step:
@@ 3.1 Gradient-Guided Candidate Selection
    A local neighborhood $\mathcal{N}_{K}(t_{c})$ is defined, consisting of the nearest neighbors $K$ to the current pivot token $t_{c}$ in the embedding space.
@@
    Within this neighborhood, the entropy of each candidate token $t\in\mathcal{N}_{K}(t_{c})$ is estimated using a first-order Taylor approximation:
    $\hat{H}(t)\approx H(t_{c})+\nabla_{e}H(t_{c})^{\top}(e_{t}-e_{t_{c}})$
    where $e_{t}$ and $e_{t_{c}}$ are the respective embedding vectors. This approximation avoids the high cost of exact entropy calculations for all neighbors.
@@
    A batch $\mathcal{B}$ of $B$ tokens with the highest approximated entropy $\hat{H}(t)$ is selected from the neighborhood. This batch is returned for verification.
  4. 4.
    To guide the next search step, the actual entropy $H_{t}$ is computed for the tokens in $\mathcal{B}$, and the one with the highest actual entropy becomes the new pivot, $t_{c}$.
Figure 2 visualizes this process. This local search strategy significantly improves the accuracy of entropy estimation by focusing on tokens close to the pivot token. This entire process corresponds to the ‘SelectCandidateBatch’ function in Algorithm 1.
@@
3.2 Robust Verification of Candidates
Once a batch of promising candidates is selected, each one undergoes a rigorous verification step (‘VerifyIsGlitch’ in our pipeline) to filter out false positives. A candidate is confirmed as a true glitch only if it fails the repetition task across a set of diverse prompt templates. To ensure this robustness, our verification process uses $m=3$ templates for cross-validation: in addition to the template used for optimization (Sec. 3.1), we employ two others adapted from prior work (Magikarp; GlitchHunter). A token is formally classified as a glitch if and only if its failure rate (as defined in Equation 1) is 1. The additional prompt templates and analysis of false positives are provided in Appendix A.1 and Appendix A.2, respectively.
3.3 Practical Implementation: Search Space Pre-filtering
To maximize search efficiency, following (land2024fishing), we perform a one-time pre-filtering of the vocabulary (‘PreFilter’ in Algorithm 1). This step removes tokens that pose no real-world risk, allowing GlitchMiner’s overall search to focus its resources. We filter three categories:
  • SPECIAL Tokens: Predefined symbols like [BOS] or </s>. These are filtered out because they serve a reserved functional role for the model (e.g., indicating the start of a sequence) rather than representing user-generated text.
  • UNDECODEABLE Tokens: Tokens that correspond to byte sequences that cannot be decoded into a valid string, often because they violate encoding standards like UTF-8. They are excluded as they do not map to any meaningful user-generated text.
  • UNREACHABLE Tokens: Tokens that exist in the vocabulary but can never be generated by the model’s tokenizer from any text input.
Further explanations of these token categories are provided in Appendix A.3.
4 Experiments
4.1 Experimental Setup
Evaluated LLMs. We used a diverse set of LLMs from five different model families to evaluate the performance of our glitch token detection approach. The selected models include Meta’s Llama series (llama2; llama3_1), Alibaba’s Qwen models (qwen2; qwen2_5), Google’s Gemma models (gemma2), Microsoft’s Phi-3 models (phi3), and Mistral models (mistral_7B; mistral2024nemo). The details are presented in Table 1.
Evaluation Metrics. We evaluate our glitch token detection method using the Detected@N metric, which counts the number of true glitch tokens identified within the top $N$ predictions. For instance, Detected@1000 measures how many glitch tokens are found among the top 1000 candidates. This metric balances detection accuracy and query efficiency, reflecting a method’s practical effectiveness under fixed query budgets. Comparing Detected@N values thus provides a direct measure of each method’s ability to maximize glitch token discovery while minimizing computational resources, making it well-suited for real-world applications.